OpenStudy (daniellelovee):

You must convert a known volume of gas from both a given temperature and a given pressure to STP. Which of the following is the best law to use?

OpenStudy (daniellelovee):

but I'll ask them separately so that you can get 5 medals

OpenStudy (daniellelovee):

@Somy

OpenStudy (somy):

what is STP? Standard temperature and pressure right?

OpenStudy (somy):

do you know the values for it?

OpenStudy (daniellelovee):

PV = nRT?

OpenStudy (somy):

hmm no it is a formula for gas law thats right but im asking about STP values there are known values of temperature, pressure and volume in STP condition

OpenStudy (daniellelovee):

Values as in numbers for the mesure of the temperature and the pressure?

OpenStudy (somy):

the starting formula is thats right PV=nRT but we are comaring P, V, and T so lets put them on one side that makes a new formula which is for ' comparison ' as this question want nR=PV/T we make nR as subject because we assume they are constant so your final formula will be \[\frac{ P_1 V_1}{ T1 } = \frac{ P_2V_2 }{ T_2 }\]

OpenStudy (somy):

nope wait i'll find it for you

OpenStudy (daniellelovee):

sorry my computer freezes at the worst moments

OpenStudy (somy):

its okay)

OpenStudy (daniellelovee):

ok ok Im here now let me see if I got this so the numbers that you wanted me to find were 273 K (0° Celsius) and the standard pressure is 1 atm pressure. At STP, one mole of gas occupies 22.4 L of volume (molar volume).

OpenStudy (daniellelovee):

is it combined gas law?

OpenStudy (daniellelovee):

@Somy ?

OpenStudy (daniellelovee):

@ichliebedich are you good at chemistry?

OpenStudy (ichliebedich):

sorta

OpenStudy (daniellelovee):

can you please check this?

OpenStudy (ichliebedich):

1. Combined gas law P1V1/T1 = P2V2/T2

OpenStudy (daniellelovee):

ok so I was right thank you ^.^

OpenStudy (ichliebedich):

no problem

OpenStudy (somy):

owh sorry i was away

OpenStudy (somy):

yep thats right you got it

OpenStudy (daniellelovee):

lol is alright thank you though it seems like I'm finally getting descent at chemistry ^.^

OpenStudy (somy):

good to know :)))